# Kolmogorov-size particles in homogeneous and isotropic turbulence

### v1

@@ 4.2 The particle-velocity Structure function
Figure 14: Second order structure function based on the particle velocity field $S_{2,p}$. Top: $\rho_{p}/\rho_{f}=5$. Bottom: $\rho_{p}/\rho_{f}=100$. The empty circles in the bottom panel are from the simulation with the coarser grid at $\Phi_{V}=10^{-3}$ and $\rho_{p}/\rho_{f}=100$, and are shown for validation purpose.
@@
The bottom panel of figure 14 deals with the $\rho_{p}/\rho_{f}=100$ cases. A first observation is that the results from the PR-DNS with $\Phi_{V}=10^{-5}$ and $\Phi_{V}=10^{-3}$ do not collapse; this is consistent with the larger flow modulation observed for the larger volume fraction, and agrees with the above discussed results for the single-particle statistics. Notably, for heavy particles $S_{2,p}$ differs from the fluid structure function $S_{2}$ at the small scales. According to both the PP-DNS and the PR-DNS, $S_{2,p}$ does not exhibit a $r^{2}$ scaling at the smallest scales, being substantially flat at small $r$. The relative motion between couples of heavy particles placed at a small distances $r$ is substantially uncorrelated as well as decoupled from the small scale fluid motion due to their large inertia. Note that, the absence of the $S_{2,p}\sim r^{2}$ scaling indicates that at small scales the Eulerian particle velocity field cannot be described with a Taylor expansion. Notably, figure 14 shows that $S_{2,p}$ recovers the $S_{2}-$slope at larger $r$, exhibiting the classical Kolmogorov $r^{2/3}$ scaling in the inertial range of scales. This suggests that, despite the large inertia, the relative particle-particle velocity $\delta\bm{u}_{p}$ between two particles is driven by turbulent eddies having size comparable to $r$, provided that $r$ is large enough. When comparing the results of the PP-DNS with those of the PR-DNS, we note that the slope of $S_{2,p}$ matches for small ($r/\eta\lessapprox 5$) and large ($r\gtrapprox 20$) scales. For intermediate scales $5\lessapprox r/\eta\lessapprox 20$, instead, the PR-DNS predict a steeper slope for both $\Phi_{V}=10^{-5}$ and $\Phi_{V}=10^{-3}$. The finite size of the particles does not influence $S_{2,p}$ for large and small scales where $r/D_{p}=\mathcal{O}(100)$ and $r/D_{p}=\mathcal{O}(1)$, but it does for intermediate scales $r/D_{p}=\mathcal{O}(10)$.
@@
Figure 15 sheds further light on the relative particle-particle velocity by plotting the distribution of $\delta\bm{u}_{p}\cdot\bm{r}/r$, i.e. the component of $\delta\bm{u}_{p}$ projected along the vector separating the two particles, for different values of $r$. When $\delta\bm{u}_{p}\cdot\bm{r}/r>0$, the two particles depart, whereas they get closer when $\delta\bm{u}_{p}\cdot\bm{r}/r<0$. We consider the case with $\rho_{p}/\rho_{f}=100$ to provide further insights of the distribution of $S_{2,p}$, shown in the right panels of figure 14. A first observation is that, similarly to what found for larger particles by Chiarini & Rosti (2024), the distribution of $\delta\bm{u}_{p}\cdot\bm{r}/r$ is left skewed, with a slightly positive mode and a long negative tail. The distribution becomes progressively more flat when increasing $r$, in agreement with a lower level of the correlation between the velocity of the two particles. When comparing the distributions for $\Phi_{V}=10^{-5}$ and $\Phi_{V}=10^{-3}$, figure 15 shows that for all $r$ the tails are shorter for the larger $\Phi_{V}$, with the difference decreasing as $r$ increases. This is consistent with the stronger flow modulation that globally leads to a weaker level of the flow fluctuations; see table 1. Also, in agreement with the evolution of $S_{2,p}$ with $r$ (see figure 14), for $\Phi_{V}=10^{-5}$ the distribution with the PP-DNS collapses nicely with that obtained with the PR-DNS at small scales (see the top panels), with some substantial differences arising for intermediate scales when the finite-size effects are relevant.
@@ 5.1 Clustering
Over the years, several tools have been used to characterise the spatial arrangement of the particles in the flow (see Brandt & Coletti, 2022, for an overview). We use the Voronoï tessellation, which has been extensively used by several authors (see for example Monchaux et al., 2010, 2012). The position of each particle is identified with its centre and the computational domain is divided in subdomains, such that each grid cell is associated with the closest particle. The Voronoï volume $V_{V}$ of each particle is thus defined as the collective volume of grid cells that are closer to it than to other particles. The inverse of the Voronoï volumes provides a measure of the local concentration: particles placed in void regions possess a large Voronoï volume, while particles that are part of a cluster have a small Voronoï volume. Based on these observations, the intensity of clustering of a suspension can be measured by comparing the distribution of its Voronoï volumes to that of a control consisting of an equivalent, uniformly random, non-overlapping suspension of particles. More intense clustering leads to a variance of the distribution of $V_{V}$ that is larger than that of the control.
@@
We now move to the effect of the volume fraction $\Phi_{V}$ and of the particle density $\rho_{p}/\rho_{f}$. As expected, the level of clustering increases with $\Phi_{V}$. When fixing $\Phi_{V}$, instead, figure 16 shows that heavier particles with $\rho_{p}/\rho_{f}=100$ cluster more than lighter particles with $\rho_{p}/\rho_{f}=5$. For light particles, the low level of clustering $\sigma/\sigma_{rand}\approx 1$ agrees with the previous results of Fiabane et al. (2012), Uhlmann & Chouippe (2017) and Chiarini & Rosti (2024), who considered larger particles $5\leq D_{p}/\eta\leq 123$ over a wide range of Reynolds numbers $105\leq Re_{\lambda}\leq 430$. Light particles have small inertia and are less likely to drift from the trajectories of the fluid elements. In contrast, the larger level of clustering observed when increasing the particle density from $\rho_{p}/\rho_{f}=5$ to $\rho_{p}/\rho_{f}=100$ is not consistent with what found for larger particles. For suspensions of particles with size in the inertial range of scales, indeed, Chiarini & Rosti (2024) found that the level of clustering exhibits a non-monotonous dependence on $\rho_{p}/\rho_{f}$, with the maximum occurring at intermediate densities (see figure 24 of their paper), and the minimum being at the largest density ratio they considered, i.e. $\rho_{p}/\rho_{f}=100$. However, they found that in the $D_{p}-\rho_{p}$ space of parameters the maximum level of clustering moves towards larger $\rho_{p}$ as $D_{p}$ decreases, suggesting that the tendency of particles to cluster is driven by the Stokes number of the particles, rather than by their density. Accordingly, their data show that the level of clustering is maximum when $St=\mathcal{O}(1-10)$. This agrees with the early works of Wang & Maxey (1993), Fessler et al. (1994) and Alised et al. (2002), and it is consistent with the our results.
@@
A different type of information regarding the spatial arrangement of the particles can be provided by means of the radial distribution function $g(r)$, also referred to as pair correlation function (see figure 18). It describes how the particle density varies as a function of the distance away from a reference particle. In other words, it is a measure of the probability of finding particles at a distance $r$ relative to that of a homogeneous distribution. Following Salazar et al. (2008) and Saw et al. (2008), the radial distribution function is defined as
@@ 5.2 Preferential sampling
In the limit where the point-particle approximation holds, Maxey (1987) has shown that particles with large density tend to collect in regions of high strain rate and low vorticity. In the presence of a vortex, indeed, heavy particles cannot follow the flow streamlines because of their large inertia, and tend to drift from the vortex core. Similarly, in the case of a pure straining flow, heavy particles drift towards the stagnation point at the centre. We quantify the tendency of particles to sample regions of high strain by using the second invariant of the deformation tensor (see §3.4), i.e. $Q=-S_{ij}S_{ij}/2+\omega^{2}/4$; recall that regions where $Q$ is large and positive are regions of high vorticity ($Q\sim\omega^{2}/4$), while regions where $Q$ is large and negative are regions of high strain ($Q\sim-S_{ij}S_{ij}$). We investigate the particle preferential sampling by computing the probability density function of $Q$ at the particles position (Squires & Eaton, 1990). For the sake of brevity, in this section we limit the investigation to the largest volume fraction $\Phi_{V}=10^{-3}$. For PP-DNS the value of $Q$ at each particle position is obtained after linear interpolation. For PR-DNS, instead, the value of $Q$ seen by each particle is estimated as the average value within a spherical shell centred with the particle and having a radius $R_{sh}>R_{p}$, where $R_{p}=D_{p}/2$ is the radius of the particles. It is important to note, however, that due to the particles’ backreaction, in the case of PR-DNS, the value of $Q$ seen by each particle is actually the result of three different effects: (i) the larger scale flow properties of the region that the particle is sampling, (ii) the smaller scale influence of the particle on the surrounding flow, and (iii) the effect of nearby particles on the flow. Also, a suitable choice of $R_{sh}$ should be made: when $R_{sh}$ is too small, only the influence of the particle on the surrounding flow is considered (see for example Kidanemariam et al., 2013; Oka & Goto, 2022), while when $R_{sh}$ is too large, spurious contributions that do not affect the particle location are instead captured. In order to obtain a more complete picture, we have tested different values of $R_{sh}$.
@@
Figure 19 shows the probability density function of $Q_{\ell}$, i.e. $Q$ computed at the particle position. We start by looking at the dependence of the PR-DNS results on the radius of the shell $R_{sh}$. For $\rho_{p}/\rho_{f}=5$, the curves computed for values of $R_{sh}$ between $2\leq R_{sh}/R_{p}\leq 7$ show an almost perfect overlap. This is consistent with the low level of modulation discussed in §3, and indicates that for these light particles the main contribution to $Q_{\ell}$ comes from the larger scale properties of the flow region sampled by the particles (note that due to the low volume fraction the influence of the particle-particle interaction is negligible). For heavy particles with $\rho_{p}/\rho_{f}=100$, instead, the distribution of $Q_{\ell}$ largely varies with $R_{sh}$: as $R_{sh}$ decreases, the left tail of the distribution becomes longer, meaning that particles are more likely to see large negative values of $Q$. These large negative values of $Q$ are, at least partially, the result of the influence of the particles on the neighbouring flow; see the $Q-R$ map in figures 8. Note that, for $R_{sh}/R_{p}\gtrapprox 5$ the distribution of $Q_{\ell}$ shows only marginal variations, as for these $R_{sh}$ the large scale flow contribution dominates. This suggests that for $\rho_{p}/\rho_{f}=100$ the influence of particles on the surrounding flow extends for less than $5R_{p}$. Overall, for both light and heavy particles the distribution of $Q_{\ell}$ is left skewed and shows an almost null probability of positive $Q_{\ell}$. At the present parameters, both PP-DNS and PR-DNS give evidence that Kolmogorov-size particles preferentially sample regions of high strain rate. This is also visible in the instantaneous snapshot shown in figure 1, with particles sampling regions with low $\omega^{2}$. In other words, in the context of Kolmogorov-size particles the formation of clusters is, at least partially, governed by the centrifuge mechanism.
@@
Let us now focus on the differences between the PR-DNS and PP-DNS results. For particles with $\rho_{p}/\rho_{f}=5$, the $Q_{\ell}$ distribution obtained with PP-DNS overlaps almost perfectly with that obtained with PR-DNS; the point-particle approximation predicts fairly well the tendency of light particles to sample the $Q<0$ regions of the flow. Note that this is consistent with the good agreement found in figure 17 when discussing the distribution of the Voronoï volumes. For $\rho_{p}/\rho_{f}=100$, instead, the $Q_{\ell}$ distribution obtained with the PP-DNS significantly deviates from that obtained with the PR-DNS. For all $R_{sh}$, the $Q_{\ell}$ distribution obtained with PR-DNS shows a shorter right tail and predicts a higher probability of negative $Q$: finite-size heavy particles are less/more likely to see positive/negative values of $Q$. Based on this, one may conclude that, at the present parameters, the PP-DNS underestimates the tendency of the particles to preferentially sample regions of high-strain, and this may explain the larger level of clustering predicted by the PR-DNS for the $\Phi_{V}=10^{-3}$ and $\rho_{p}/\rho_{f}=100$ case (see figure 16).
@@
A last comment regards the influence of $\rho_{p}/\rho_{f}$ on the distribution of $Q_{\ell}$. According to both PP-DNS and PR-DNS, heavier particles exhibit a somewhat larger tendency to sample regions with more negative $Q$, as shown by the left tail being longer for the $\rho_{p}/\rho_{f}=100$ case. Due to their larger inertia, indeed, heavier particles enhance the centrifuge mechanism, being more likely to drift from the high-vorticity regions of the flow.

### v2

@@ 4.2 The particle-velocity Structure function
Figure 15: Second order structure function based on the particle velocity field $S_{2,p}$. Top: $\rho_{p}/\rho_{f}=5$. Bottom: $\rho_{p}/\rho_{f}=100$. The empty circles in the bottom panel are from the simulation with the coarser grid at $\Phi_{V}=10^{-3}$ and $\rho_{p}/\rho_{f}=100$, and are shown for validation purpose. The inset in the bottom panel shows the compensated $S_{2,p}r^{-2/3}$ in the $10\leq r/\eta\leq 50$ range.
@@
The bottom panel of figure 15 deals with the $\rho_{p}/\rho_{f}=100$ cases. A first observation is that the results from the PR-DNS with $\Phi_{V}=10^{-5}$ and $\Phi_{V}=10^{-3}$ do not collapse; this is consistent with the larger flow modulation observed for the larger volume fraction, and agrees with the above discussed results for the single-particle statistics. Notably, for heavy particles $S_{2,p}$ differs from the fluid structure function $S_{2}$ at the small scales. According to both the PP-DNS and the PR-DNS, $S_{2,p}$ does not exhibit a $r^{2}$ scaling at the smallest scales, being substantially flat at small $r$. The relative motion between couples of heavy particles placed at a small distances $r$ is substantially uncorrelated as well as decoupled from the small scale fluid motion due to their large inertia. Note that, the absence of the $S_{2,p}\sim r^{2}$ scaling indicates that at small scales the Eulerian particle velocity field cannot be described with a Taylor expansion. Notably, figure 15 shows that $S_{2,p}$ recovers the $S_{2}-$slope at larger $r$, exhibiting the classical Kolmogorov $r^{2/3}$ scaling in the inertial range of scales. This suggests that, despite the large inertia, the relative particle-particle velocity $\delta\bm{u}_{p}$ between two particles is driven by turbulent eddies having size comparable to $r$, provided that $r$ is large enough. When comparing the results of the PP-DNS with those of the PR-DNS, we note that the slope of $S_{2,p}$ matches for small ($r/\eta\lessapprox 5$) and large ($r\gtrapprox 30$) scales. For intermediate scales $10\lessapprox r/\eta\lessapprox 30$, instead, the PR-DNS predict a steeper slope for both $\Phi_{V}=10^{-5}$ and $\Phi_{V}=10^{-3}$ (see the inset in the bottom panel of figure 15). The finite size of the particles does not influence $S_{2,p}$ for large and small scales where $r/D_{p}=\mathcal{O}(100)$ and $r/D_{p}=\mathcal{O}(1)$, but it does for intermediate scales $r/D_{p}=\mathcal{O}(10)$.
@@
Figure 16 sheds further light on the relative particle-particle velocity by plotting the distribution of $\delta\bm{u}_{p}\cdot\bm{r}/r$, i.e. the component of $\delta\bm{u}_{p}$ projected along the vector separating the two particles, for different values of $r$. When $\delta\bm{u}_{p}\cdot\bm{r}/r>0$, the two particles depart, whereas they get closer when $\delta\bm{u}_{p}\cdot\bm{r}/r<0$. We consider the case with $\rho_{p}/\rho_{f}=100$ to provide further insights of the distribution of $S_{2,p}$, shown in the bottom panel of figure 15. A first observation is that, similarly to what found for larger particles by Chiarini & Rosti (2024), the distribution of $\delta\bm{u}_{p}\cdot\bm{r}/r$ is left skewed, with a slightly positive mode and a long negative tail. The distribution becomes progressively more flat when increasing $r$, in agreement with a lower level of the correlation between the velocity of the two particles. When comparing the distributions for $\Phi_{V}=10^{-5}$ and $\Phi_{V}=10^{-3}$, figure 16 shows that for all $r$ the tails are shorter for the larger $\Phi_{V}$, with the difference decreasing as $r$ increases. This is consistent with the stronger flow modulation that globally leads to a weaker level of the flow fluctuations; see table 1. Also, in agreement with the evolution of $S_{2,p}$ with $r$ (see figure 15), for $\Phi_{V}=10^{-5}$ the distribution obtained with the PP-DNS collapses nicely with that obtained with the PR-DNS at small scales (see the top panels), with some substantial differences arising for intermediate scales when the finite-size effects are relevant.
@@ 5.1 Clustering
Over the years, several tools have been used to characterise the spatial arrangement of the particles in the flow (see Brandt & Coletti, 2022, for an overview). We use the Voronoï tessellation, which has been extensively used by several authors (see for example Monchaux et al., 2010, 2012). The position of each particle is identified with its centre and the computational domain is divided in subdomains, such that each grid cell is associated with the closest particle. The Voronoï volume $V_{V}$ of each particle is thus defined as the collective volume of grid cells that are closer to it than to other particles. The inverse of the Voronoï volumes provides a measure of the local concentration: particles placed in void regions possess a large Voronoï volume, while particles that are part of a cluster have a small Voronoï volume. Based on these observations, the intensity of the clustering of a suspension can be measured by comparing the distribution of its Voronoï volumes to that of a control consisting of an equivalent, uniformly random suspension of particles. More intense clustering leads to a variance of the distribution of $V_{V}$ that is larger than that of the control. In the context of PR-DNS, the overlap between particles is not allowed also in the reference random arrangement.
@@
We now move to the effect of the volume fraction $\Phi_{V}$ and of the particle density $\rho_{p}/\rho_{f}$. As expected, the level of clustering increases with $\Phi_{V}$. When fixing $\Phi_{V}$, instead, figure 17 shows that heavier particles with $\rho_{p}/\rho_{f}=100$ cluster more than lighter particles with $\rho_{p}/\rho_{f}=5$. For light particles, the low level of clustering $\sigma/\sigma_{rand}\approx 1$ agrees with the previous results of Fiabane et al. (2012), Uhlmann & Chouippe (2017) and Chiarini & Rosti (2024), who considered larger particles $5\leq D_{p}/\eta\leq 123$ over a wide range of Reynolds numbers $105\leq Re_{\lambda}\leq 430$. Light particles have small inertia and are less likely to drift from the trajectories of the fluid elements. In contrast, the larger level of clustering observed when increasing the particle density from $\rho_{p}/\rho_{f}=5$ to $\rho_{p}/\rho_{f}=100$ is not consistent with what found for larger particles. For suspensions of particles with size in the inertial range of scales, indeed, Chiarini & Rosti (2024) found that the level of clustering exhibits a non-monotonous dependence on $\rho_{p}/\rho_{f}$, with the maximum occurring at intermediate densities (see figure 24 of their paper), and the minimum being at the largest density ratio they considered, i.e. $\rho_{p}/\rho_{f}=100$. However, they found that in the $D_{p}-\rho_{p}$ space of parameters the maximum level of clustering moves towards larger $\rho_{p}$ as $D_{p}$ decreases, suggesting that the tendency of particles to cluster is driven by the Stokes number of the particles, rather than by their density. Accordingly, their data show that the level of clustering is maximum when $St=\mathcal{O}(1-10)$. This agrees with the early works of Wang & Maxey (1993), Fessler et al. (1994) and Aliseda et al. (2002), and it is consistent with our results (see table 1 for the particles’ Stokes number).
@@
A different type of information regarding the spatial arrangement of the particles can be provided by means of the radial distribution function $g(r)$, also referred to as pair correlation function; see figure 19. It describes how the particles’ density varies as a function of the distance away from a reference particle. In other words, it is a measure of the probability of finding particles at a distance $r$ relative to that of a homogeneous distribution. Following Salazar et al. (2008) and Saw et al. (2008), the radial distribution function is defined as
@@ 5.2 Preferential sampling
In the limit where the point-particle approximation holds, Maxey (1987) has shown that particles with large $\rho_{p}/\rho_{f}$ tend to collect in regions of high strain rate and low vorticity. In the presence of a vortex, indeed, heavy particles cannot follow the flow streamlines because of their large inertia, and tend to drift from the vortex core. Similarly, in the case of a pure straining flow, heavy particles drift towards the stagnation point at the centre. We quantify the tendency of particles to sample regions of high strain by using the second invariant of the deformation tensor (see §3.4), i.e. $Q=-S_{ij}S_{ij}/2+\omega^{2}/4$; recall that regions where $Q$ is large and positive are regions of high vorticity ($Q\sim\omega^{2}/4$), while regions where $Q$ is large and negative are regions of high strain ($Q\sim-S_{ij}S_{ij}$). We investigate the particle preferential sampling by computing the probability density function of $Q$ at the particles position (Squires & Eaton, 1990). For the sake of brevity, in this section we limit the investigation to the largest volume fraction $\Phi_{V}=10^{-3}$. For PP-DNS the value of $Q$ at each particle position is obtained after linear interpolation. For PR-DNS, instead, the value of $Q$ seen by each particle is estimated as the average value within a spherical shell centred with the particle and having a radius $R_{sh}>R_{p}$, where $R_{p}=D_{p}/2$ is the radius of the particles. It is important to note, however, that due to the particles’ backreaction, in the case of PR-DNS, the value of $Q$ seen by each particle is actually the result of three different effects: (i) the larger scale flow properties of the region that the particle is sampling, (ii) the smaller scale influence of the particle on the surrounding flow, and (iii) the effect of nearby particles on the flow. Also, a suitable choice of $R_{sh}$ should be made: when $R_{sh}$ is too small, only the influence of the particle on the surrounding flow is considered (see for example Kidanemariam et al., 2013; Oka & Goto, 2022), while when $R_{sh}$ is too large, spurious contributions that do not affect the particle location are instead captured. In order to obtain a complete picture, we have tested different values of $R_{sh}$.
@@
Figure 20 shows the probability density function of $Q_{\ell}$, i.e. $Q$ computed at the particle position. We start by looking at the dependence of the PR-DNS results on the radius $R_{sh}$ of the shell. For $\rho_{p}/\rho_{f}=5$, the curves computed for values of $R_{sh}$ between $2\leq R_{sh}/R_{p}\leq 7$ show an almost perfect overlap. This is consistent with the low level of modulation discussed in §3, and indicates that for these light particles the main contribution to $Q_{\ell}$ comes from the larger scale properties of the flow region sampled by the particles (note that due to the low volume fraction the influence of the particle-particle interaction is negligible). For heavy particles with $\rho_{p}/\rho_{f}=100$, instead, the distribution of $Q_{\ell}$ largely varies with $R_{sh}$: as $R_{sh}$ decreases, the left tail of the distribution becomes longer, meaning that particles are more likely to see large negative values of $Q$. These large negative values of $Q$ are, at least partially, the result of the influence of the particles on the neighbouring flow; see the $Q-R$ map in figures 8. Note that, for $R_{sh}/R_{p}\gtrapprox 5$ the distribution of $Q_{\ell}$ shows only marginal variations, as for these $R_{sh}$ the large scale flow contribution dominates. This suggests that for $\rho_{p}/\rho_{f}=100$ the influence of particles on the surrounding flow extends for less than $5R_{p}$. Overall, for both light and heavy particles the distribution of $Q_{\ell}$ is left skewed and shows an almost null probability of positive $Q_{\ell}$. At the present parameters, both PP-DNS and PR-DNS give evidence that Kolmogorov-size particles preferentially sample regions of high strain rate. This is also visible in the instantaneous snapshot shown in figure 1, with particles sampling regions with low $\omega^{2}$. In other words, in the context of Kolmogorov-size particles the formation of clusters is, at least partially, governed by the centrifuge mechanism.
@@
Let us now focus on the differences between the PR-DNS and PP-DNS results. For particles with $\rho_{p}/\rho_{f}=5$, the $Q_{\ell}$ distribution obtained with PP-DNS overlaps almost perfectly with that obtained with PR-DNS; the point-particle approximation predicts fairly well the tendency of light particles to sample the $Q<0$ regions of the flow. Note that this is consistent with the good agreement found in figure 18 when discussing the distribution of the Voronoï volumes. For $\rho_{p}/\rho_{f}=100$, instead, the $Q_{\ell}$ distribution obtained with the PP-DNS significantly deviates from that obtained with the PR-DNS. For all $R_{sh}$, the $Q_{\ell}$ distribution obtained with PR-DNS shows a shorter right tail and predicts a higher probability of negative $Q$: finite-size heavy particles are less/more likely to see positive/negative values of $Q$. Based on this, one may conclude that, at the present parameters, the PP-DNS underestimates the tendency of the particles to preferentially sample regions of high-strain, and this may explain the larger level of clustering predicted by the PR-DNS for the $\Phi_{V}=10^{-3}$ and $\rho_{p}/\rho_{f}=100$ case (see figure 17). We speculate that the discrepancy between the PR-DNS and the PP-DNS observed for the $\Phi_{V}=10^{-3}$ and $\rho_{p}/\rho_{f}=100$ case is due to the flow modulation, rather than to the particles’ finite-size effects. Unlike for the light particles ($\rho_{p}/\rho_{f}=5$), indeed, at the present parameters the influence of the heavy particles ($\rho_{p}/\rho_{f}=100$) on the fluid phase is not negligible (see §3).
@@
A last comment regards the influence of $\rho_{p}/\rho_{f}$ on the distribution of $Q_{\ell}$. According to both PP-DNS and PR-DNS, heavier particles exhibit a larger tendency to sample regions with more negative $Q$, as shown by the left tail being longer for the $\rho_{p}/\rho_{f}=100$ case. Due to their larger inertia, indeed, heavier particles enhance the centrifuge mechanism, being more likely to drift from the high-vorticity regions of the flow.
[…]